OpenStudy (anonymous):

if f(x)=7x+11 and g(x) = x^2-5x-3 write an equation for the case where the outputs of the two functions are equal.

OpenStudy (anonymous):

f(x) = g(x)\[7x+11=x ^{2}-5x-3\]\[x ^{2}-12x-14=0\]A solution may then be found by use of the quadratic formula.